\begin{thebibliography}{99}                                                                                               %


\bibitem {Bol98}B. Bollob\'{a}s, \emph{Modern Graph Theory}\textit{,} Graduate
Texts in Mathematics, 184, Springer-Verlag, New York (1998).

\bibitem {BoNi04}B. Bollob\'{a}s, V. Nikiforov, Joints in graphs, to appear in
\emph{Discrete Math.}

\bibitem {BoNi07}B. Bollob\'{a}s, V. Nikiforov, Cliques and the spectral
radius, \emph{J. Combin. Theory Ser. B.} \textbf{97} (2007), 859-865.

\bibitem {Erd68}P. Erd\H{o}s, On some new inequalities concerning extremal
properties of graphs, in: \emph{Theory of Graphs (Proc. Colloq., Tihany,
1966),} pp. 77--81, Academic Press, New York, 1968.

\bibitem {Nik07}V. Nikiforov, Graphs with many $r$-cliques have large complete
$r$-partite subgraphs, to appear in \emph{Bull. London Math. Soc.} Update
available at \emph{http://arxiv.org/math.CO/0703554}

\bibitem {Nik07a}V. Nikiforov, Stability for large forbidden graphs, submitted
for publication. Preprint available at \emph{http://arxiv.org/abs/0707.2563}

\bibitem {Nik02}V. Nikiforov, Some inequalities for the largest eigenvalue of
a graph, \emph{Combin. Probab. Comp.} \textbf{11} (2002), 179-189.

\bibitem {Nik06}V. Nikiforov, The smallest eigenvalue of $K_{r}$-free graphs,
\emph{Discrete Math.} \textbf{306} (2006), 612-616.

\bibitem {Nik07c}V. Nikiforov, Eigenvalues and forbidden subgraphs I,
\emph{Linear Algebra Appl.} \textbf{422} (2007), 384-390.

\bibitem {Nik07d}V. Nikiforov, Bounds on graph eigenvalues II, \emph{Linear
Algebra Appl.} \textbf{427} (2007) 183-189.

\bibitem {Nik07e}V. Nikiforov, A spectral condition for odd cycles, to appear
in \emph{Linear Algebra Appl. }Update available at
\emph{http://arxiv.org/abs/0707.4499}

\bibitem {Nik07f}V. Nikiforov, More spectral bounds on the clique and
independence numbers, submitted for publication. Preprint available at
\emph{http://arxiv.org/abs/0706.0548}

\bibitem {Nik07g}V. Nikiforov, A spectral Erd\H{o}s-Stone-Bollob\'{a}s
theorem, submitted for publication. Preprint available at
\emph{http://arxiv.org/abs/0707.2259}

\bibitem {Nik07h}V. Nikiforov, Spectral saturation: inverting the spectral
Tur\'{a}n theorem, submitted for publication. Preprint available at
\emph{http://arxiv.org/abs/0707.2259}

\bibitem {Sim68}M. Simonovits, A method for solving extremal problems in graph
theory, stability problems, in: \emph{Theory of Graphs (Proc. Colloq., Tihany,
1966),} pp. 279--319, Academic Press, New York, 1968.
\end{thebibliography}

\documentclass[12pt]{article}%
\usepackage{graphicx}
\usepackage[intlimits]{amsmath}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{amssymb}%
 \makeatletter
\newfont{\footsc}{cmcsc10 at 8truept}
\newfont{\footbf}{cmbx10 at 8truept}
\newfont{\footrm}{cmr10 at 10truept}
\newtheorem{theorem}{Theorem}

\newtheorem{fact}[theorem]{Fact}

\newenvironment{proof}[1][Proof]{\noindent{\textbf {#1}  }}  {\hfill$\Box$\bigskip}

\begin{document}

\title{A spectral stability theorem for large forbidden subgraphs}
\author{Vladimir Nikiforov\\{\small Department of Mathematical Sciences, University of Memphis, Memphis,
TN 38152}\\{\small email: vnikifrv@memphis.edu}}
\maketitle

\begin{abstract}
Let $\mu\left(  G\right)  $ be the largest eigenvalue of a graph $G,$ let
$K_{r}\left(  s_{1},\ldots,s_{r}\right)  $ be the complete $r$-partite graph
with parts of size $s_{1},\ldots,s_{r},$ and let $T_{r}\left(  n\right)  $ be
the $r$-partite Tur\'{a}n graph of order $n.$ Our main result is:\medskip

For all $r\geq2$ and all sufficiently small $c>0,$ $\varepsilon>0,$ every
graph $G$ of sufficiently large order $n$ with $\mu\left(  G\right)
\geq\left(  1-1/r-\varepsilon\right)  n$ satisfies one of the conditions:

(a) $G$ contains a $K_{r+1}\left(  \left\lfloor c\ln n\right\rfloor
,\ldots,\left\lfloor c\ln n\right\rfloor ,\left\lceil n^{1-\sqrt{c}%
}\right\rceil \right)  ;$

(b) $G$ differs from $T_{r}\left(  n\right)  $ in fewer than $\left(
\varepsilon^{1/4}+c^{1/\left(  8r+8\right)  }\right)  n^{2}$ edges.\medskip

In particular, this result strengthens the stability theorem of Erd\H{o}s and
Simonovits.\medskip

\textbf{Keywords: }\textit{stability, forbidden subgraphs, }$r$%
\textit{-partite subgraphs; largest eigenvalue of a graph; spectral Tur\'{a}n
theorem.}

\end{abstract}

This note is part of an ongoing project aiming to build extremal graph theory
on spectral grounds, see, e.g., \cite{BoNi07} and \cite{Nik02,Nik07h}.

Let $\mu\left(  G\right)  $ be the largest adjacency eigenvalue of a graph
$G,$ let $K_{r}\left(  s_{1},\ldots,s_{r}\right)  $ be the complete
$r$-partite graph with parts of size $s_{1},\ldots,s_{r},$ and let
$T_{r}\left(  n\right)  $ be the $r$-partite Tur\'{a}n graph of order $n$. In
\cite{Nik07a} we extended the Erd\H{o}s-Simonovits stability theorem
\cite{Erd68}, \cite{Sim68} as:\medskip

\emph{Let }$r\geq2,$\emph{ }$1/\ln n<c<r^{-3\left(  r+14\right)  \left(
r+1\right)  },$ $0<\varepsilon<r^{-24},$\emph{ and }$G$\emph{ be a graph of
order }$n.$\emph{ If }$G$\emph{ has }$\left\lceil \left(  1-1/r-\varepsilon
\right)  n^{2}/2\right\rceil $\emph{ edges, then }$G$\emph{ satisfies one of
the conditions:}

\emph{(a) }$G$\emph{ contains a }$K_{r+1}\left(  \left\lfloor c\ln
n\right\rfloor ,\ldots,\left\lfloor c\ln n\right\rfloor ,\left\lceil
n^{1-\sqrt{c}}\right\rceil \right)  ;$

\emph{(b) }$G$\emph{ differs from }$T_{r}\left(  n\right)  $\emph{ in fewer
than }$\left(  \varepsilon^{1/3}+c^{1/\left(  3r+3\right)  }\right)  n^{2}%
$\emph{ edges.}\medskip

Here we derive essentially the same conclusion from the weaker premise
$\mu\left(  G\right)  >\left(  1-1/r-\varepsilon\right)  n:$

\begin{theorem}
\label{th1}Let $r\geq2,$ $1/\ln n<c<r^{-8\left(  r+21\right)  \left(
r+1\right)  },$ $0<\varepsilon<2^{-36}r^{-24},$ and $G$ be a graph of order
$n.$ If $\mu\left(  G\right)  >\left(  1-1/r-\varepsilon\right)  n$, then $G$
satisfies one of the conditions:

(a) $G$ contains a $K_{r+1}\left(  \left\lfloor c\ln n\right\rfloor
,\ldots,\left\lfloor c\ln n\right\rfloor ,\left\lceil n^{1-\sqrt{c}%
}\right\rceil \right)  ;$

(b) $G$ differs from $T_{r}\left(  n\right)  $ in fewer than $\left(
\varepsilon^{1/4}+c^{1/\left(  8r+8\right)  }\right)  n^{2}$ edges.
\end{theorem}

\subsubsection*{Remarks}

\begin{itemize}
\item[-] Since $\mu\left(  G\right)  $ is at least the average degree of $G,$
Theorem \ref{th1} implies essentially the above extension of the
Erd\H{o}s-Simonovits stability theorem.

\item[-] The relation between $c$ and $n$ in Theorem \ref{th1} needs
explanation. First, for fixed $c,$ it shows how large must be $n$ to get a
valid conclusion. But, in fact, the relation is subtler, for $c$ itself may
depend on $n,$ e.g., letting $c=1/\ln\ln n,$ the conclusion is meaningful for
sufficiently large $n.$

\item[-] Choosing randomly a graph of order $n$ with $\left\lceil \left(
1-1/r\right)  n^{2}/2\right\rceil $ edges, we can find a graph containing no
$K_{2}\left(  \left\lfloor c^{\prime}\ln n\right\rfloor ,\left\lfloor
c^{\prime}\ln n\right\rfloor \right)  $ and differing from $T_{r}\left(
n\right)  $ in more that $c^{\prime\prime}n^{2}$ edges for some positive
$c^{\prime}$ and $c^{\prime\prime},$ independent of $n$. Hence, condition
\emph{(a)} is essentially best possible.

\item[-] The factor $\varepsilon^{1/4}+c^{1/\left(  8r+8\right)  }$ in
condition \emph{(b)} is far from the best one, but is simple.\bigskip
\end{itemize}

To prove Theorem \ref{th1}, we introduce two supporting results. Our notation
follows \cite{Bol98}; given a graph $G,$ we write:\medskip

- $\left\vert G\right\vert $ for the number of vertices set of $G;$

- $e\left(  G\right)  $ for the number of edges of $G;$

- $\delta\left(  G\right)  $ for the minimum degree of $G;$

- $k_{r}\left(  G\right)  $ for the number of $r$-cliques of $G.$\bigskip

An\emph{ }$r$\emph{-joint }of size $t$ is the union of $t$ distinct
$r$-cliques sharing an edge. We write $js_{r}\left(  G\right)  $ for the
maximum size of an $r$-joint in a graph $G.$\bigskip

The following two facts play crucial roles in our proof.

\begin{fact}
[\cite{Nik07h}, Theorem 4]\label{stabj} Let $r\geq2,$ $0<b<2^{-10}r^{-6},$
$n\geq r^{20},$ and $G$ be a graph of order $n.$ If $\mu\left(  G\right)
>\left(  1-1/r-b\right)  n,$ then $G$ satsisfies one of the conditions:

(i) $js_{r+1}\left(  G\right)  >n^{r-1}/r^{2r+5};$

(ii) $G$ contains an induced $r$-partite subgraph $G_{0}$ satisfying
$\left\vert G_{0}\right\vert \geq\left(  1-4b^{1/3}\right)  n$ and
$\delta\left(  G_{0}\right)  >\left(  1-1/r-7b^{1/3}\right)  n.$
\end{fact}

\begin{fact}
[\cite{Nik07}, Theorem 1]\label{ES}Let $r\geq2,$ $c^{r}\ln n\geq1,$ and $G$ be
a graph of order $n$. If $k_{r}\left(  G\right)  \geq cn^{r},$ then $G$
contains a $K_{r}\left(  s,\ldots,s,t\right)  $ with $s=\left\lfloor c^{r}\ln
n\right\rfloor $ and $t>n^{1-c^{r-1}}.\hfill\square$
\end{fact}

\bigskip

\begin{proof}
[\textbf{Proof of Theorem \ref{th1}}]Let $G$ be a graph of order $n$ with
$\mu\left(  G\right)  >\left(  1-1/r-\varepsilon\right)  n.$ Define the
procedure $\mathcal{P}$ as follows:\medskip

\textbf{While}\emph{ }$js_{r+1}\left(  G\right)  >n^{r-1}/r^{2r+5}$
\textbf{do}

\qquad\emph{Select an edge contained in }$\left\lceil n^{r-1}/r^{2r+5}%
\right\rceil $\emph{ cliques of order }$r+1$ \emph{and remove it from }%
$G.$\medskip

Set for short $\theta=c^{1/\left(  r+1\right)  }r^{2r+5}$ and assume first
that $\mathcal{P}$ removes at least $\left\lceil \theta n^{2}\right\rceil $
edges before stopping. Then
\[
k_{r+1}\left(  G\right)  \geq\theta n^{r-1}/r^{2r+5}=c^{1/\left(  r+1\right)
}n^{r+1},
\]
and Fact \ref{ES} implies that $K_{r+1}\left(  \left\lfloor c\ln
n\right\rfloor ,\ldots,\left\lfloor c\ln n\right\rfloor ,\left\lceil
n^{1-\sqrt{c}}\right\rceil \right)  \subset G.$ Thus condition \emph{(a)}
holds, completing the proof.

Assume now that $\mathcal{P}$ removes fewer than $\left\lceil \theta
n^{2}\right\rceil $ edges before stopping; write $G^{\prime}$ for the
resulting graph.

Letting $\mu\left(  X\right)  $ be the largest eigenvalue of a Hermitian
matrix $X,$ recall Weyl's inequality
\[
\mu\left(  B\right)  \geq\mu\left(  A\right)  -\mu\left(  A-B\right)  ,
\]
holding for any Hermitian matrices $A$ and $B.$ Also, recall that $\mu\left(
H\right)  \leq\sqrt{2e\left(  H\right)  }$ for any graph $H.$ Applying these
results to the graphs $G$ and $G^{\prime},$ we find that
\[
\mu\left(  G^{\prime}\right)  \geq\mu\left(  G\right)  -\sqrt{2\theta}%
n\geq\left(  1-1/r-\varepsilon-\sqrt{2\theta}\right)  n.
\]

From $\ln n\geq1/c\geq r^{8\left(  r+21\right)  \left(  r+1\right)  }$ we
easily get $n>r^{20}.$ Set for short $a=\left(  \varepsilon+\sqrt{2\theta
}\right)  ^{1/3}.$ Since
\[
\varepsilon+\sqrt{2\theta}\leq2^{-36}r^{-24}+2r^{-4\left(  r+21\right)
\left(  r+1\right)  }<2^{-10}r^{-6},
\]
and $js_{r+1}\left(  G^{\prime}\right)  \leq n^{r-1}/r^{2r+5},$ Fact
\ref{stabj} implies that $G^{\prime}$ contains an induced $r$-partite subgraph
$G_{0},$ satisfying $\left\vert G_{0}\right\vert \geq\left(  1-4a\right)  n$
and $\delta\left(  G_{0}\right)  >\left(  1-1/r-7a\right)  n.$

Let $V_{1},\ldots,V_{r}$ be the parts of $G_{0}.$ For every $i\in\left[
r\right]  ,$ we see that%
\[
\left\vert V_{i}\right\vert \geq n-%
%TCIMACRO{\tsum \limits_{s\in\left[  r\right]  \backslash\left\{  i\right\}
%}}%
%BeginExpansion
{\textstyle\sum\limits_{s\in\left[  r\right]  \backslash\left\{  i\right\}  }}
%EndExpansion
\left\vert V_{s}\right\vert \geq n-\left(  r-1\right)  \left(  n-\delta\left(
G_{0}\right)  \right)  \geq\left(  1/r-7\left(  r-1\right)  a\right)  n.
\]
For each $i\in\left[  r\right]  ,$ select a set $U_{i}\subset V_{i}$ with
\[
\left\vert U_{i}\right\vert =\left\lceil \left(  1/r-7\left(  r-1\right)
a\right)  n\right\rceil ,
\]
and write $G_{1}$ for the graph induced by $\cup_{i=1}^{r}U_{i}$. Clearly
$G_{1}$ can be made complete $r$-partite by adding at most
\[
\left(  \left(  1-1/r\right)  \left\vert G_{1}\right\vert -\delta\left(
G_{1}\right)  \right)  \left\vert G_{1}\right\vert /2
\]
edges. We see that%
\[
\delta\left(  G_{1}\right)  \geq\delta\left(  G_{0}\right)  -\left\vert
G_{0}\right\vert +\left\vert G_{1}\right\vert \geq-n/r-4an+\left\vert
G_{1}\right\vert ,
\]
and so,
\begin{align*}
\left(  1-1/r\right)  \left\vert G_{1}\right\vert -\delta\left(  G_{1}\right)
&  \leq\left(  1/r+4a\right)  n-\left\vert G_{1}\right\vert /r\\
&  =\left(  1/r+4a\right)  n-\left(  1/r-7\left(  r-1\right)  a\right)
n=7ran.
\end{align*}
Therefore, $G_{1}$ can be made complete $r$-partite by adding at most
\[
7ran\left\vert G_{1}\right\vert /2<4ran^{2}%
\]
edges.

The complete $r$-partite graph with parts $U_{1},\ldots,U_{r}$ can be
transformed into $T_{r}\left(  n\right)  $ by changing at most $\left(
n-\left\vert G_{1}\right\vert \right)  n$ edges. Since
\[
\left(  n-\left\vert G_{1}\right\vert \right)  n\leq\left(  n-r\left(
1/r-7\left(  r-1\right)  a\right)  n\right)  n=7r\left(  r-1\right)  an^{2},
\]
we find that $G$ differs from $T_{r}\left(  n\right)  $ in at most $\left(
\theta+\left(  7r^{2}-3r\right)  a\right)  n^{2}$ edges. Now, condition
\emph{(b)} follows in view of
\begin{align*}
\theta+\left(  7r^{2}-3r\right)  a  &  =\theta+\left(  7r^{2}-3r\right)
\left(  \varepsilon+\sqrt{2\theta}\right)  ^{1/3}<\theta+8r^{2}\varepsilon
^{1/3}+2\left(  7r^{2}-3r\right)  \theta^{1/6}\\
&  <8r^{2}\varepsilon^{1/3}+r^{6}\theta^{1/6}<\varepsilon^{1/4}+r^{\left(
2r+5\right)  /6+6}c^{1/\left(  6r+6\right)  }\leq\varepsilon^{1/4}%
+c^{1/\left(  8r+8\right)  }.
\end{align*}
The proof is completed.
\end{proof}

\subsubsection*{Concluding remark}

Finally, a word about the project mentioned in the introduction: in this
project we aim to give wide-range results that can be used further, adding
more integrity to spectral extremal graph theory.

\end{document}